# Non-Hermitian Origin of Wannier Localizability and Detachable Topological Boundary States

### v2

@@ IV Intrinsic non-Hermitian topology
$\tilde{H}\left(\bm{k}\right)\coloneqq\begin{pmatrix}0&H\left(\bm{k}\right)\\ H^{{\dagger}}\left(\bm{k}\right)&0\end{pmatrix}_{\sigma},$ (IV.1)
which respects chiral symmetry $\Gamma H\left(\bm{k}\right)\Gamma^{-1}=-H\left(\bm{k}\right)$ with $\Gamma=\sigma_{z}$ by construction. When the non-Hermitian Hamiltonian $H\left(\bm{k}\right)$ has a point gap, the Hermitized Hamiltonian $\tilde{H}\left(\bm{k}\right)$ also has a gap, and vice versa. Thus, the topological classification of $H\left(\bm{k}\right)$ coincides with that of $\tilde{H}\left(\bm{k}\right)$, which is denoted by $K_{\rm P}$. By contrast, the presence of a line gap imposes an additional constraint on $\tilde{H}\left(\bm{k}\right)$. In fact, a non-Hermitian Hamiltonian $H\left(\bm{k}\right)$ with a real or an imaginary line gap can be continuously deformed to a Hermitian or an anti-Hermitian Hamiltonian while keeping the line gap and relevant symmetry, respectively [41]. Notably, Hermiticity or anti-Hermiticity of $H\left(\bm{k}\right)$ imposes another chiral symmetry with $\Gamma_{\rm r}=\sigma_{y}$ or $\Gamma_{\rm i}=\sigma_{x}$ on the Hermitized Hamiltonian $\tilde{H}\left(\bm{k}\right)$, respectively. This additional chiral symmetry leads to the different topological classification denoted by $K_{\rm L_{r}}$ or $K_{\rm L_{i}}$ in the presence of a real or an imaginary line gap (and a point gap as well). Forgetting $\Gamma_{\rm r}$ ($\Gamma_{\rm i}$) defines a homomorphism $f_{\rm r}:K_{\rm L_{r}}\to K_{\rm P}$ ($f_{\rm i}:K_{\rm L_{i}}\to K_{\rm P}$) from $K_{\rm L_{r}}$ ($K_{\rm L_{i}}$) to $K_{\rm P}$. Owing to the dimensional isomorphism of the $K$-theory, it is sufficient to compute $f_{\rm r}$ or $f_{\rm i}$ in zero dimension to obtain those in arbitrary dimensions. Table S1 summarizes these homomorphisms in the Altland-Zirnbauer symmetry classes, which is relevant to this work. See Refs. [51, 92] for the homomorphisms in all the 38 symmetry classes.
@@
$\begin{array}[]{ccccccccccccc}\hline\cr\hline\cr\mbox{Class}&\mbox{Gap}&d=0&d=1&d=2&d=3&d=4&d=5&d=6&d=7\\ \hline\cr{\rm A}&{\rm L}\to{\rm P}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}\\ &&&&&&&&&\\ \hline\cr{\rm AIII}&{\rm L_{r}}\to{\rm P}&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}&{\mathbb{Z}}\to 0\\ &&&&&&&&&\\ &{\rm L_{i}}\to{\rm P}&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0\\ &&(n,m)\mapsto n-m&&(n,m)\mapsto n-m&&(n,m)\mapsto n-m&&(n,m)\mapsto n-m&\\ \hline\cr{\rm AI^{\dagger}}&{\rm L}\to{\rm P}&{\mathbb{Z}}\to 0&0\to 0&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}\\ &&&&&&&&n\mapsto 0&\\ \hline\cr{\rm BDI^{\dagger}}&{\rm L_{r}}\to{\rm P}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\\ &&&&&&&&&n\mapsto 0\\ &{\rm L_{i}}\to{\rm P}&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&0\to 0&0\to 0&2{\mathbb{Z}}\oplus 2{\mathbb{Z}}\to 2{\mathbb{Z}}&0\to 0&{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\\ &&(n,m)\mapsto n-m&&&&(n,m)\mapsto n-m&&(n,m)\mapsto n+m&(n,m)\mapsto n+m\\ \hline\cr{\rm D^{\dagger}}&{\rm L_{r}}\to{\rm P}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}\\ &&n\mapsto 0&&&&&&&\\ &{\rm L_{i}}\to{\rm P}&{\mathbb{Z}}\to{\mathbb{Z}}_{2}&0\to{\mathbb{Z}}&0\to 0&0\to 0&2{\mathbb{Z}}\to 0&0\to 2{\mathbb{Z}}&{\mathbb{Z}}_{2}\to 0&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\\ &&n\mapsto n&&&&&&&n\mapsto n\\ \hline\cr{\rm DIII^{\dagger}}&{\rm L_{r}}\to{\rm P}&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0\\ &&&n\mapsto 0&&&&&&\\ &{\rm L_{i}}\to{\rm P}&{\mathbb{Z}}\to{\mathbb{Z}}_{2}&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&{\mathbb{Z}}\to 0&0\to 0&{\mathbb{Z}}\to 2{\mathbb{Z}}&0\to 0\\ &&n\mapsto n&&n\mapsto 2n&&&&n\mapsto n&\\ \hline\cr{\rm AII^{\dagger}}&{\rm L}\to{\rm P}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0&0\to 0&0\to 2{\mathbb{Z}}\\ &&&&n\mapsto 0&&&&&\\ \hline\cr{\rm CII^{\dagger}}&{\rm L_{r}}\to{\rm P}&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0&0\to 0\\ &&&&&n\mapsto 0&&&&\\ &{\rm L_{i}}\to{\rm P}&2{\mathbb{Z}}\oplus 2{\mathbb{Z}}\to 2{\mathbb{Z}}&0\to 0&{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}\oplus{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0&0\to 0&0\to 0\\ &&(n,m)\mapsto n-m&&(n,m)\mapsto n+m&(n,m)\mapsto n+m&(n,m)\mapsto n-m&&\\ \hline\cr{\rm C^{\dagger}}&{\rm L_{r}}\to{\rm P}&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0&0\to 0\\ &&&&&&n\mapsto 0&&&\\ &{\rm L_{i}}\to{\rm P}&2{\mathbb{Z}}\to 0&0\to 2{\mathbb{Z}}&{\mathbb{Z}}_{2}\to 0&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}\to{\mathbb{Z}}_{2}&0\to{\mathbb{Z}}&0\to 0&0\to 0\\ &&&&&n\mapsto n&n\mapsto n&&&\\ \hline\cr{\rm CI^{\dagger}}&{\rm L_{r}}\to{\rm P}&0\to 0&0\to 0&0\to 2{\mathbb{Z}}&2{\mathbb{Z}}\to 0&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}&{\mathbb{Z}}_{2}\to{\mathbb{Z}}&{\mathbb{Z}}\to 0\\ &&&&&&&n\mapsto 0&&\\ &{\rm L_{i}}\to{\rm P}&{\mathbb{Z}}\to 0&0\to 0&{\mathbb{Z}}\to 2{\mathbb{Z}}&0\to 0&{\mathbb{Z}}\to{\mathbb{Z}}_{2}&0\to{\mathbb{Z}}_{2}&{\mathbb{Z}}\to{\mathbb{Z}}&0\to 0\\ &&&&n\mapsto n&&n\mapsto n&&n\mapsto 2n&\\ \hline\cr\hline\cr\end{array}$
V Reflection-invariant topological crystalline insulators
Our formulation is also applicable to topological crystalline insulators. As a prime example, we study a two-dimensional topological insulator with commutative chiral and reflection symmetries [108, 109, 110]:
$\Gamma\mathcal{H}\left(k_{x}\right)\Gamma^{-1}=-\mathcal{H}\left(k_{x}\right),\quad\mathcal{R}_{+}\mathcal{H}\left(k_{x}\right)\mathcal{R}_{+}^{-1}=\mathcal{H}\left(-k_{x}\right),$ (V.1)
@@ V Reflection-invariant topological crystalline insulators
$\Gamma^{2}=\mathcal{R}_{+}^{2}=1,\quad\left[\Gamma,\mathcal{R}_{+}\right]=0.$ (V.2)
At a reflection-invariant edge, a gapless state
$\mathcal{H}_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)=k_{x}\sigma_{x}$ (V.3)
appears. This Dirac model in Eq. (I.5) indeed respects these symmetries with $\Gamma=\sigma_{z}$ and $\mathcal{R}_{+}=\sigma_{z}$.
Notably, this reflection-invariant gapless state exhibits no spectral flow and hence is detachable from the bulk, similarly to the Dirac surface state protected by chiral symmetry. To confirm this, we couple it to a trivial band $\varepsilon\sigma_{x}$ ($\varepsilon\geq 0$) by
~$\tilde{\mathcal{H}}_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)=\begin{pmatrix}k_{x}\sigma_{x}&v\sigma_{-}\\ v\sigma_{+}&\varepsilon\sigma_{x}\end{pmatrix},$ (V.4)
with the coupling strength $v\geq 0$. This four-band model remains to respect chiral symmetry with $\Gamma=\sigma_{z}$ and reflection symmetry with $\mathcal{R}_{+}=\mathrm{diag}\left(\sigma_{z},1\right)$. Similarly to the chiral-symmetric two-dimensional Dirac model studied in Sec. I.3, the four energy bands are obtained as
$\sqrt{2}E\left(k_{x}\right)=\pm\sqrt{k^{2}_{x}+\varepsilon^{2}+v^{2}\pm\sqrt{\left(k^{2}_{x}-\varepsilon^{2}+v^{2}\right)^{2}+4\varepsilon^{2}v^{2}}},$ (V.5)
@@
$\Gamma H_{{\rm AIII}+\mathcal{R}_{+}}^{{\dagger}}\left(k_{x}\right)\Gamma^{-1}=-H_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right),\quad\mathcal{R}_{+}H_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)\mathcal{R}_{+}^{-1}=H_{{\rm AIII}+\mathcal{R}_{+}}\left(-k_{x}\right),\quad\left[\Gamma,\mathcal{R}_{+}\right]=0$ (V.6)

### v3

@@ V Intrinsic non-Hermitian topology
$\tilde{H}\left(\bm{k}\right)\coloneqq\begin{pmatrix}0&H\left(\bm{k}\right)\\ H^{{\dagger}}\left(\bm{k}\right)&0\end{pmatrix}_{\sigma},$ (V.1)
which respects chiral symmetry $\Gamma H\left(\bm{k}\right)\Gamma^{-1}=-H\left(\bm{k}\right)$ with $\Gamma=\sigma_{z}$ by construction. When the non-Hermitian Hamiltonian $H\left(\bm{k}\right)$ has a point gap, the Hermitized Hamiltonian $\tilde{H}\left(\bm{k}\right)$ also has a gap, and vice versa. Thus, the topological classification of $H\left(\bm{k}\right)$ coincides with that of $\tilde{H}\left(\bm{k}\right)$, which is denoted by $K_{\rm P}$. By contrast, the presence of a line gap imposes an additional constraint on $\tilde{H}\left(\bm{k}\right)$. A non-Hermitian Hamiltonian $H\left(\bm{k}\right)$ with a real or an imaginary line gap can be continuously deformed to a Hermitian or an anti-Hermitian Hamiltonian while keeping the line gap and relevant symmetry, respectively [41]. Hermiticity or anti-Hermiticity of $H\left(\bm{k}\right)$ imposes another chiral symmetry with $\Gamma_{\rm r}=\sigma_{y}$ or $\Gamma_{\rm i}=\sigma_{x}$ on the Hermitized Hamiltonian $\tilde{H}\left(\bm{k}\right)$, respectively. This additional chiral symmetry leads to the different topological classification denoted by $K_{\rm L_{r}}$ or $K_{\rm L_{i}}$ in the presence of a real or an imaginary line gap (and a point gap as well). Forgetting $\Gamma_{\rm r}$ ($\Gamma_{\rm i}$) defines a homomorphism $f_{\rm r}:K_{\rm L_{r}}\to K_{\rm P}$ ($f_{\rm i}:K_{\rm L_{i}}\to K_{\rm P}$) from $K_{\rm L_{r}}$ ($K_{\rm L_{i}}$) to $K_{\rm P}$. Owing to the dimensional isomorphism of the $K$-theory, it is sufficient to compute $f_{\rm r}$ or $f_{\rm i}$ in zero dimension to obtain those in arbitrary dimensions. Table S1 summarizes these homomorphisms in the Altland-Zirnbauer symmetry classes, which is relevant to this work. See Refs. [51, 92] for the homomorphisms in all the 38 symmetry classes.
@@
VI Reflection-invariant topological crystalline insulators
Our formulation is also applicable to topological crystalline insulators. As a prime example, we study two-dimensional topological insulators with commutative chiral and reflection symmetries [108, 109, 110]:
$\Gamma\mathcal{H}\left(k_{x},k_{y}\right)\Gamma^{-1}=-\mathcal{H}\left(k_{x},k_{y}\right),\quad\mathcal{R}_{+}\mathcal{H}\left(k_{x},k_{y}\right)\mathcal{R}_{+}^{-1}=\mathcal{H}\left(-k_{x},k_{y}\right),$ (VI.1)
@@ VI Reflection-invariant topological crystalline insulators
$\Gamma^{2}=\mathcal{R}_{+}^{2}=1,\quad\left[\Gamma,\mathcal{R}_{+}\right]=0.$ (VI.2)
To apply our formalism, we focus on effective gapless Hamiltonians at boundaries. At an edge around a reflection-invariant corner, a gapless state
$\mathcal{H}_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)=k_{x}\sigma_{x}$ (VI.3)
appears, where $k_{x}$ is a momentum along the edge of the two-dimensional system. This Dirac model indeed respects the symmetries with $\Gamma=\sigma_{z}$ and $\mathcal{R}_{+}=\sigma_{z}$. Notably, this reflection-invariant gapless state exhibits no spectral flow and hence is detachable from the bulk, similarly to the Dirac surface state protected by chiral symmetry. To confirm this, we couple it to a trivial band $\varepsilon\sigma_{x}$ ($\varepsilon\geq 0$) by
$\tilde{\mathcal{H}}_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)=\begin{pmatrix}k_{x}\sigma_{x}&v\sigma_{+}\\ v\sigma_{-}&\varepsilon\sigma_{x}\end{pmatrix}$, (VI.4)
with the coupling strength $v\geq 0$. This four-band model still respects chiral symmetry with $\Gamma=\sigma_{z}$ and reflection symmetry with $\mathcal{R}_{+}=\mathrm{diag}\left(\sigma_{z},1\right)$. Similarly to the chiral-symmetric two-dimensional Dirac model studied in Sec. I.3, the four energy bands are obtained as
$\sqrt{2}E\left(k_{x}\right)=\pm\sqrt{k^{2}_{x}+\varepsilon^{2}+v^{2}\pm\sqrt{\left(k^{2}_{x}-\varepsilon^{2}+v^{2}\right)^{2}+4\varepsilon^{2}v^{2}}},$ (VI.5)
@@
$\Gamma H_{{\rm AIII}+\mathcal{R}_{+}}^{{\dagger}}\left(k_{x}\right)\Gamma^{-1}=-H_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right),\quad\mathcal{R}_{+}H_{{\rm AIII}+\mathcal{R}_{+}}\left(k_{x}\right)\mathcal{R}_{+}^{-1}=H_{{\rm AIII}+\mathcal{R}_{+}}\left(-k_{x}\right),\quad\left[\Gamma,\mathcal{R}_{+}\right]=0.$ (VI.6)